OpenStudy (pumpkin1102):

Josh and his friends bought vanilla wafers for $4 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $45 to buy a total of 27 packets of wafers of the two varieties.

OpenStudy (pumpkin1102):

Part A: Write a system of equations that can be solved to find the number of packets of vanilla wafers and the number of packets of chocolate wafers that Josh and his friends bought at the carnival. Define the variables used in the equations. Part B: How many packets of chocolate wafers and vanilla wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer.

OpenStudy (pumpkin1102):

plz help I have literally been stuck on this question for 1 week and a half

OpenStudy (mertsj):

How many packets did they buy in all?

OpenStudy (thereisnousername):

a = vanilla b = chocolate a + b = 27 packets 4a + 1b = $45 so now you have a simultaneous equation, can you solve from here?

OpenStudy (pumpkin1102):

change the positive B to negative so that equation would become a = 27 - b so that would be 4a + (27 - b) = 45 am I right so far? @ThereisnoUsername

OpenStudy (thereisnousername):

yes, now what do you do?

OpenStudy (pumpkin1102):

Multiply 4 x 27 and get 108 and then I am stuck again

OpenStudy (pumpkin1102):

@ThereisnoUsername